OpenStudy (anonymous):

A country's population in 1990 was 132 million. In 1998, 137 million. Estimate the population in 2003. (exponential growth, P=Ae^kt, round k four decimal places)

OpenStudy (anonymous):

Plug-in for t=0 to solve for A: 132 = Ae^k(0) A = 132 in 1998, t=1998-1990=8 years and population P=137 million, plug-in values to solve for k: 137 = 132e^8k 137/132 = e^8k = 1.038 Applying ln on both sides to cancel e: 8k = ln 1.038 k = 0.0046 Estimating population in 2003: t = 2003 - 1990 = 13 years Therefore P(2003) = 132e^(0.0046*13) P(2003) = approx 140.13 million // answer